introduction to Short-Term Financial Management 1 27 Total Liabilities Common Stock Retained Earnings $1.151.00 $1110.00 $120.00 $676.80 $796.80 120.00 330.00 450.00 $1,560 tal Shareholders Equity Total Liabilities & shareholders Equity $ 1,947.80 .00 2. From the financial statements shown in problem 1, calculate the AFNwc. From the financial statements showniroblem 1, calculate g ffor year T+1). Assuming revenues of $5,050 in year t, how does between the years T and T+1? this value compare to the realized growth in revenues Use the following incomplete cash flow statement to solve for the change in cash. TATEMENT OF CASH FLOWS perations $144.00 $15.00 $400.00 $400.00 reciation ounts Receivable
PROBLEMS income statement and balance sheet shown statement J Use the cash flows. Gross Profit Depreclation Interest Expense the finan $231.20 From the fina revenues of s between the Use the fol STATLMUNT Balance Sheet YEAR T+1 + Deprec $714.80 $500.00 300.00 $1,51480 $660.00 $227.00 $433.00 $1,94780 Cash and Equivalents Accounts Receivable Inventory Total Current Assets 34000 Fixed Assets $800.00 Accumulated Depreciation Net Fixed Assets Total Assets Accounts Payable Short-Term Notes Payable $320.00 40.00 $11.00 $371.00 780.00 Total Current Liabilities Long-Term Debt $360.00 750.00
introduction to Short-Term Financial Manac $1110.00 Total Liabilities Common Stock 51,151.00 $120.00 $676.80 $796.80 5120.00 $330.00 $450.00 Aetained Earnings tal Shareholders Equity tal Liabilities & Shareholders Equity $ 1,947.80 $1,560.00 From the financial statements shown in problem 1, calculate the AFNWC. om the financial statements shown igProblem 1, calculate g for year T+1) roblem 1, calculate g T+1 enues of $5,050 in year t, how does this value compare to the realized growth veen the years T and T+1? ke following incomplete cash flow statement to solve for the change in cast NT OF CASH FLOWS ons $144.00 $15.00 $400.00 $400.00 $140.00 $1000 on Receivable yable
Answer #1

AFNWC = Additional finance needed for working capital

= AReceivables + AInventory - AAccount Payables

Where Delta stands for change.

Based on the financial statements,

AReceivablesReceivablest+1 -Receivablest

= 500 - 500 = 0

Inventory-Inventoryt+1 _ Inventor!

=300 - 340 = - 40

Account Payables-Account Payable st+1-Account Payable st

= 320 - 300 = 20

Hence, AFNWC = 0 = (-40) - 20 = - $ 60

Hence, Net working capital will get released to the extent of $ 60.

------------------

Next question

Calculation of g*

g* stands for sustainable growth rate and is given by

gx =

Where,

NPM (net profit margin) = NI/Revenues = 346.80 / 5,700 = 6.08%

DPO (dividend payout ratio) = Dividends / NI = 0

D/E (debt-to-equity ratio) = Total Liabilities/Equity = 1,151 / 796.80 = 1.44

A/R (assets-to-revenues ratio) = Total Assets/Revenues = 1,947.8 / 5,700 = 0.34

Hence

0.0608 × (1-0) × (1+ 1.44) 0.34-0.0608x (1-0) × (1 1.44)

= 77.07%

Actual growth in Revenue = RT+1 / RT - 1 = 5,700 / 5,050 -1 = 12.87%

Actual growth in the year T+1 is far lower than the sustainable growth rate. This might be because the firm didn't raise external capital up to its full potential to drive growth. Since external capital infused were lower, the growth rate achieved was lower.